OpenStudy (anonymous):

Using graph paper, graph the following equation. Then click on the graph until the correct one is displayed. 3x - y - 1 = 0

DivineSolar (divinesolar):

Ok what do we do first to this equation?

DivineSolar (divinesolar):

Before you guys give out the answer i would suggest looking at that.

DivineSolar (divinesolar):

Answermyquestions made a good point in the post i was in.

DivineSolar (divinesolar):

MichelleiSMine is not trying i see.

OpenStudy (anonymous):

You need to get it into the form y=mx+b so just rearrange the equation it gave you 3x-y-1=0 -y-1=-3x -y=-3x+1 y=3x-1

DivineSolar (divinesolar):

cooco_crazy.